bills coffee shop makes a blend that is a mixture of two types of coffee. Type a coffee costs bill $5.70 per pound and type b coffee costs $4.65 per pound. This month bill made 181 pounds of the blend, for a total cost of $943.50. How many pounds of type A coffee did he use?

Answer by stanbon(75887)

bills coffee shop makes a blend that is a mixture of two types of coffee. Type a coffee costs bill $5.70 per pound and type b coffee costs $4.65 per pound. This month bill made 181 pounds of the blend, for a total cost of $943.50. How many pounds of type A coffee did he use?
--------
Equation:
value + value = value
5.7a + 4.65b = 943.50
a + b = 181
-------------------------
Substitute for "b" and solve for "a"::
5.7a + 4.65(181-a) = 943.50
-----
5.7a + 841.65 - 4.65a = 943.50
------
101.85/1.05 = a
a = 97 lbs (amt. of A used in the blend)
-------
Cheers,
Stan H.
--------